OpenStudy (anonymous):

Complete the equation for the linear function whose graph contains the points (9, 7) and (4, –8). y −____ =____(x − 9)

OpenStudy (freckles):

this is point slope form

OpenStudy (freckles):

they entered in for x1 the value 9 so we know y1 should be what?

OpenStudy (freckles):

\[y-y_1=m(x-x_1) \\ y-y_1=m(x-9)\]

OpenStudy (anonymous):

9

OpenStudy (freckles):

when x is 9 y is?

OpenStudy (freckles):

look at your ordered pairs given

OpenStudy (anonymous):

7

OpenStudy (freckles):

you have (9,7) and (4,-8) (4,-8) tells us when x=4 y=-8 so (9,7) tells us when x=9 y=?

OpenStudy (freckles):

yes

OpenStudy (freckles):

so y1 =7 since x1=9

OpenStudy (freckles):

\[y-y_1=m(x-x_1) \\ y-y_1=m(x-9) \\ y-7=m(x-9)\]

OpenStudy (freckles):

we talking about how to find the slope on one of your questions

OpenStudy (freckles):

remember this: \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (freckles):

use the two points given to find the slope

OpenStudy (anonymous):

ok give second to figure it out plz

OpenStudy (anonymous):

-15/-5

OpenStudy (anonymous):

if i simplify it i would get -3/-1 and then get -3

OpenStudy (anonymous):

so it would be y-7=-3(x-9)?

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

\[m=\frac{7-(-8)}{9-4}=\frac{15}{5}=3\] -3/-1 shouldn't be -3

OpenStudy (freckles):

should be 3

OpenStudy (freckles):

-/-=+

OpenStudy (freckles):

yes?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

oh i was seeing if you where there so it would be y-7=3(x-9)?

OpenStudy (freckles):

yes the slope is 3 and that line is correct

OpenStudy (anonymous):

thanks :)